How many grams of H2SO4 are produced from 32.8 grams of NaOH?


1
Expert's answer
2021-04-11T23:46:52-0400

Equation for the reaction will be

H2SO4+2NaOH→Na2SO4+2H20

Mol ratio 1:2:1:2

Number of moles of NaOH used=32.8/40=0.82 moles

This means number of moles of H2SO4 used=0.82/2=0.41moles

and Na2SO4=0.41 moles

H2SO4 Used in grams =98grams/mol x0.41 moles=40.18 grams

Na2SO4 produced in grams=142g/molx0.41moles=58.22 grams